The height of the body after 5 s from the ground is (g=9.8m/s2).

Grade:12th pass

2 Answers

Pawan Prajapati
2 years ago
Our expert is working on this class X maths answer. We will update the answer very soon.
Yash Patari
2 years ago
DEAR STUDENT
The height of the body after 5 s from the ground is (g=9.8m/s2)?
ANSWER

v=u+gt
t=5 sec
u=0
v=5g=5×9.8=49 m/s

THANKS
TEAM ASKIITIANS